OpenStudy (anonymous):

15 radical 5 is radical 225, right? @jkim3592

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@Zale101

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

15 radical 5 is 15 radical 5 u cant simplify anymore

OpenStudy (anonymous):

No I need it to EQUAL radical 225

OpenStudy (anonymous):

oh 15 radical 5 has to equal 225?

OpenStudy (anonymous):

225 equals 15 times 15

OpenStudy (anonymous):

so leave 15 alone and you ahve to do something with radical 5 to be 15

OpenStudy (anonymous):

\[\sqrt{5}\times something has \to equal 15\]

OpenStudy (anonymous):

3? I'm confused.

OpenStudy (anonymous):

radical 25 equals 5

OpenStudy (anonymous):

as in example

OpenStudy (anonymous):

so your job is to make radical 225 to make 15

OpenStudy (anonymous):

which is radical 5 times radical 45

OpenStudy (anonymous):

so the answer is radical 45

OpenStudy (mertsj):

\[15\sqrt{5}=\sqrt{15^2(5)}=\sqrt{1125}\]

OpenStudy (anonymous):

\[15\sqrt{5}\times \sqrt{45}=225\]

OpenStudy (anonymous):

My choices to pick from are 15 25 15 radical 5 10 radical 2

OpenStudy (anonymous):

wats the question?

OpenStudy (anonymous):

The question is "What number equals radical 225?"

OpenStudy (anonymous):

oh its 15 than

OpenStudy (anonymous):

because radical 225 = 15times 15 which is 15

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

\[\sqrt{225}=\sqrt{15}\times \sqrt{15}\]